OpenStudy (cavazos_899579):

Do I put these three into an equation to equal 180 to find the congruent sides? In Triangle ABC, m

OpenStudy (bibby):

yeah, first add the angles up because you know the interior angles add up to 180 then you know that the sides opposite the equal angles are congruent

OpenStudy (cavazos_899579):

Okay So i did the problem and the answers are angles A and C because their measures are 65* and x = 15.

OpenStudy (bibby):

 \(4x + 5+ 2x + 20 +5x - 10=180\\11x+15=180\\11x=165\\x=15\) plug that back into the triangle

OpenStudy (bibby):



OpenStudy (bibby):

m<A= 4x + 5=65 m<B = 2x + 20 = 50 m<C = 5x - 10. 65 the sides opposite the equal angles are confruent

OpenStudy (bibby):

\(CB\cong AB\)